You have a 3.00-liter container filled with N₂ at 25°C and 2.15 atm pressure connected to a 2.00-liter container filled with Ar at 25°C and 2.75 atm pressure. A stopcock connecting the containers is opened and the gases are allowed to equilibrate between the two containers. What is the final pressure in the two containers if the temperature remains at 25°C? Assume ideal behavior.

( Answer is asked to be in atm)
Answer #1

Applying PV = constant for both nitrogen and argon .

For nitrogen:

3.00L × 2.15atm = 5.00L × P1

P1 = 3.00×2.15/5.00 = 1.29atm

For Argon :

2.00L × 2.75atm = 5.00L × P2

P2 = 2.00×2.75/5.00 = 1.10atm

Total pressure = P1 + P2 = 1.29 + 1.10 = 2.39 atm (Answer)